OpenStudy (anonymous):

What are the steps to find the square root of 9/16?

OpenStudy (shamim):

\[\sqrt{\frac{ 9 }{ 16 }}=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

How do I figure it out though?

OpenStudy (anonymous):

How did you find it's 3/4

OpenStudy (anonymous):

\[ \sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\] In this case a = 9 and b = 16.